OpenStudy (anonymous):

Evaluate the expression (3/7)^−3.

OpenStudy (anonymous):



TheSmartOne (thesmartone):

\[(\frac{ a }{ b })^(-3)=(\frac{ b }{ a })^3\]

OpenStudy (anonymous):

7/3^3

TheSmartOne (thesmartone):

no it will be (7/3)^3

OpenStudy (anonymous):

oh...ok

TheSmartOne (thesmartone):

which will be (7^3)/(3^3)

TheSmartOne (thesmartone):

7^3=7x7x7=343 3^3=3x3x3=27

OpenStudy (anonymous):

343/27

TheSmartOne (thesmartone):

Exactly...

OpenStudy (anonymous):

thanksss :DD

OpenStudy (anonymous):

\[\frac{1}{\left(\frac{3}{7}\right)^3}=\frac{343}{27} \]

OpenStudy (anonymous):

hahhaah thank you too!